OpenStudy (studytheworld):

Help!! WIll medal and fan!! Pleaseee help Suppose an isosceles triangle ABC has A = 45 (degrees) and b=c=3. What is the length of a^2?

OpenStudy (studytheworld):

@ParthKohli @zepdrix Anyone?

OpenStudy (anonymous):

a=BC^2=9.42

OpenStudy (studytheworld):

Dude that's not even an answer choice, stop telling me false information

OpenStudy (studytheworld):

@Peach-es1234

OpenStudy (anonymous):

What is the answer choices.............

OpenStudy (ronlover101):

sin A/(BC) =sin C/(4) BC=sin A/sin C *4= 0.707*4/0.92=3.07 standard ratio of sin to side

OpenStudy (ronlover101):

so a=BC^2=9.42

OpenStudy (studytheworld):

2.30 5.27 30.73 12.73 Do one of those say 9.42? No

OpenStudy (anonymous):

Well how are we suppose to know if that was correct if the answer choices weren't posted.

OpenStudy (trojanpoem):

 use the cosine law: a^2 = b^2 + c^2 - 2bc cos A well, it's asking for a^2 , and you have all the unknowns a^2 = (3)^2 + (3)^2 - 2 * 3 * 3 * cos45 a^2 = 5.27

OpenStudy (studytheworld):

@TrojanPoem Yay thank yuo you helped me al ot

OpenStudy (trojanpoem):

Any time.

OpenStudy (trojanpoem):

@ronlover101 , Why ?

OpenStudy (ronlover101):

I was joking you rock!